OpenStudy (anonymous):

Solve each of the following problems using a system of equations. Show all of your work and state your solution in a complete sentence. 1.Tony and Belinda have a combined age of 56. Belinda is 8 more than twice Tony’s age. How old is each? (2 points) 2.Salisbury High School decided to take their students on a field trip to a theme park. A total of 150 people went on the trip. Adults pay $45.00 for a ticket and students pay $28.50 for a ticket. How many students and how many adults went to the park if they paid a total of $4770? (2 points) 3.Your piggy bank has a total of 46 coins in it; somSolve each of the following problems using a system of equations. Show all of your work and state your solution in a complete sentence. 1.Tony and Belinda have a combined age of 56. Belinda is 8 more than twice Tony’s age. How old is each? (2 points) 2.Salisbury High School decided to take their students on a field trip to a theme park. A total of 150 people went on the trip. Adults pay $45.00 for a ticket and students pay $28.50 for a ticket. How many students and how many adults went to the park if they paid a total of $4770? (2 points) 3.Your piggy bank has a total of 46 coins in it; som

OpenStudy (anonymous):

some are dimes and some are quarters. If you have a total of $7.00, how many quarters and how many dimes do you have? (2 points) 4.Follow the 5 steps below to complete this problem. (4 points) Step 1: Pick a friend or family member to be the character of your word problem. This friend or family member may do one of the following: •Drive a boat •Drive a jet ski Step 2: Select a current speed of the water in mph. Step 3: Select the number of hours (be reasonable please) that your friend or family member drove the boat or jets ski against the current speed you chose in step 2. Step 4: Select the number of hours that your friend or family member made the same trip with the current (this should be a smaller number, as your friend or family member will be traveling with the current). Step 5: Write out the word problem you created and calculate how fast your friend or family member was traveling in still water. Round your answer to the nearest mph.

OpenStudy (anonymous):

. Let x = Tony's age then from "Belinda is 8 more than twice Tony’s age." 2x+8 = Belinda's age . From:"Tony and Belinda have a combined age of 56" we get our equation: x + 2x+8 = 56 3x+8 = 56 3x = 48 x = 16 years old (Tony) . Belinda: 2x+8 = 2(16)+8 = 16+8 = 24 years old

OpenStudy (anonymous):

24 years